OpenStudy (anonymous):

how do i solve for x? 3^(2x-1) = 1/9

sam (.sam.):

\[3^{2x-1} = 1/9\] \[\Large 3^{2x-1} = 9^{-1}\] \[\Large 3^{2x-1} = 3^{-2}\] 2x-1=-2 x=?

OpenStudy (anonymous):

thanks